A ring of radius 15 cm is injected with an amount of charge of 3.5 × 10−12C, uniformly distributed along the ring. The ring rotates with an angular velocity of 17 rad/s. Find the power radiated by the system.
Answer #1

Power is given by,

P = V*I

here, V = Electric potential due to ring = k*q/r

I = q/t

given, q = charge = 3.5*10^-12 C

k = 9*10^9

r = radius = 15 cm = 0.15 m

t = time period = 2*pi/w

w = angular speed = 17 rad/s

So,

P = V*I = (k*q/r)*(q*w/(2*pi))

Using known values:

P = (9*10^9)*((3.5*10^-12)^2)*17/(2*pi*0.15)

P = 2.0*10^-12 W